OpenStudy (anonymous):

Write an equation of a line parallel to line GH below in slope-intercept form that passes through the point (-5, 6). y = −x − 1 y = x − 1 y = −x + 11 y = x + 11

OpenStudy (anonymous):

a line through the points G. (-2,2) and H. (2,6)

OpenStudy (nikato):

can you find the slope of GH?

OpenStudy (anonymous):

I got 8?

OpenStudy (nikato):

not quite. remember how to find your slope? \[\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\]

OpenStudy (nikato):

and plug in your points. \[\frac{ 6-2 }{ 2-(-2) }\]

OpenStudy (anonymous):

4/4

OpenStudy (nikato):

correct. and you can just simplify to just 1. and what do you know about the slopes of two parallel lines?

OpenStudy (triciaal):

parallel lines have the same slope eqn of a line in point slope format is y-y1= m(x-x1) where you have 2 points on the line (x,y) and the given point (x1,y1)

OpenStudy (anonymous):

They have the same slope.

OpenStudy (anonymous):

I got y=x-1?

OpenStudy (nikato):

well... i got something different. since you know parallel lines have the same slope, the other line also has a slope of 1. and you know it has the point( -5,6) plug it into the point-sloep form to find the y- intercept.

OpenStudy (nikato):